Maria is three times as old as Joel. In 6 years, Maria will be twice as old as Joel. How old is Maria now?
Joe is x years old.
Maria is 3x.
In 6 years time, Joe is (x+6)
Maria is 2(x+6)=2x+12
So, it should be
3x+6=2(x+6)
3x+6=2x+12
3x-2x=12-6
x=6
Now, to check this
3(6)+6=2(6+6)
18+6=2(12)
24=24

Answer by Einstein99(1)

Maria is three times as old as Joel. In 6 years, Maria will be twice as old as Joel. How old is Maria now?
Joe is x years old.
Maria is 3x.
In 6 years time, Joe is (x+6)
Maria is 2(x+6)=2x+12
So, it should be
3x+6=2(x+6)
3x+6=2x+12
3x-2x=12-6
x=6
Now, to check this
3(6)+6=2(6+6)
18+6=2(12)
24=24